OpenStudy (anonymous):

Find the value of the sum. n integral i=1 (3+4i)^2

OpenStudy (anonymous):

summoning @sirm3d :D

OpenStudy (sirm3d):

can you type the using the equation button?

OpenStudy (sirm3d):

or maybe draw

OpenStudy (anonymous):

Ok one second.

OpenStudy (anonymous):

They really should make an app for iPad...

OpenStudy (anonymous):

\[\sum_{i=1}^{n}(3+4i)^2\]

OpenStudy (sirm3d):

we'll need three formulas \[\sum_{i=1}^{n}k=kn\] \[\sum_{i=1}^{n}i=\frac{ n(n+1) }{ 2 }\] \[\sum_{i=1}^{n}i^2 =\frac{ n(n+1)(2n+1) }{ 6 }\]

OpenStudy (sirm3d):

just expand the squared binomial, then apply the formula above

OpenStudy (anonymous):

9+ 24i + 16 i^2 and then sub i and i^2 into those formulas?

OpenStudy (sirm3d):

yes

OpenStudy (anonymous):

i'm no done yet, but is it suppose to be like tis?

OpenStudy (sirm3d):

it should be (n+1), not (n-1)

OpenStudy (anonymous):

oh..memorized it wrong.. ok..doing it again

OpenStudy (sirm3d):

just change the minus sign in your computation

OpenStudy (anonymous):

yea, i did. is everything else correct? just need to combine like terms?

OpenStudy (sirm3d):

yep. and don't miss the term n^3

OpenStudy (anonymous):

16/3 n^3 + 20n^2 + 29n :)

OpenStudy (sirm3d):

got it right

OpenStudy (anonymous):

thanks :DD

OpenStudy (anonymous):

@sirm3d when i plug this in wolfralpha, how come it's like this ? http://www.wolframalpha.com/widgets/view.jsp?id=5fd7cbbc45010f147c06926c44aff0b7

OpenStudy (anonymous):

the answer i got is 16/3 n^3 + 20n^2 + 71/3 n

OpenStudy (sirm3d):

integral is for continuous variables, summation is for discrete variables

OpenStudy (anonymous):

oohh. ok got it :) ty

OpenStudy (anonymous):

then there is no way to check my answer :/

OpenStudy (sirm3d):

here's a fourth formula \[\sum_{i=1}^{n}i^3=n^2(n+1)^2/4\]

OpenStudy (anonymous):

but there is no i^3 in my function

OpenStudy (sirm3d):

is this a summation problem or an integral problem?

OpenStudy (sirm3d):

the link gave the equation 4x^3-2x+2